# EW D-optimal Designs for Experiments with Mixed Factors

### v1

@@ 2.2 Characteristics of EW D-optimal designs
In this section, to characterize EW D-optimal designs, we fix either a prior distribution $Q(\cdot)$ on $\boldsymbol{\Theta}$ for integral-based EW D-optimality or a set of bootstrapped or sampled parameter vectors $\{\hat{\boldsymbol{\theta}}_{1},\ldots,\hat{\boldsymbol{\theta}}_{B}\}$ for sample-based EW D-optimality, so that either $E\left\{{\mathbf{F}}({\mathbf{x}}_{i},\boldsymbol{\Theta})\right\}$ is defined as in (2), or $\hat{E}\left\{{\mathbf{F}}({\mathbf{x}}_{i},\boldsymbol{\Theta})\right\}$ is defined as in (4).
For sample-based EW D-optimal designs under general parametric models, we need the regularity conditions (A1), (A2), and (B3) described in Section 2.4 of Fedorov and Leonov (2014). For integral-based EW D-optimal designs defined as in (1), we need an additional assumption (A3) (see Section S2 in the Supplementary Material).
Theorem 2.1.
For sample-based EW D-optimality under Assumptions (A1), (A2), and (B3), or integral-based EW D-optimality under Assumptions (A1), (A3), and (B3),
  • (i)
    there exists an optimal design $\boldsymbol{\xi}^{*}$ that contains no more than $p(p+1)/2$ design points;
  • (ii)
    the set of optimal designs is convex;
  • (iii)
    a design $\boldsymbol{\xi}$ is EW D-optimal if and only if $\max_{\mathbf{x}\in\mathcal{X}}d(\mathbf{x},\boldsymbol{\xi})\leq p$, where
    $d(\mathbf{x},\boldsymbol{\xi})=\left\{\begin{array}[]{cl}{\rm tr}([E\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\Theta})\}]^{-1}E\{{\mathbf{F}}({\mathbf{x}},\boldsymbol{\Theta})\}),&\mbox{for integral-based EW D-optimality;}\\ {\rm tr}([\hat{E}\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\Theta})\}]^{-1}\hat{E}\{{\mathbf{F}}({\mathbf{x}},\boldsymbol{\Theta})\}),&\mbox{for sample-based EW D-optimality.}\end{array}\right.$
@@
The proof of Theorem 2.1, as well as proofs of other theorems in this paper, is relegated to Section S4 in the Supplementary Material.
@@ 3 Algorithms for EW D-optimal Designs
3.1 Lift-one algorithm for EW D-optimal designs with discrete factors only
When all factors are discrete, that is, each $I_{j}$ is a finite set of numerical levels and the design region ${\cal X}=\prod_{j=1}^{d}I_{j}$ contains only a finite number of distinct experimental settings. In this case, we may denote $m=\prod_{j=1}^{d}|I_{j}|$ and the corresponding design settings as ${\mathbf{x}}_{1},\ldots,{\mathbf{x}}_{m}$ The EW D-optimal design problem in this case is to find ${\mathbf{w}}^{*}=(w_{1}^{*},\ldots,w_{m}^{*})^{T}\in S=\{(w_{1},\ldots,w_{m})^{T}\in\mathbb{R}^{m}\mid w_{i}\geq 0,i=1,\ldots,m;\sum_{i=1}^{m}w_{i}=1\}$, which maximizes $f_{\rm EW}({\mathbf{w}})=|\sum_{i=1}^{m}w_{i}E\{{\mathbf{F}}({\mathbf{x}}_{i},\boldsymbol{\Theta})\}|$ or $f_{\rm SEW}({\mathbf{w}})=|\sum_{i=1}^{m}w_{i}\hat{E}\{{\mathbf{F}}({\mathbf{x}}_{i},$ $\boldsymbol{\Theta})\}|$.
The original lift-one algorithm under a general setup (see Algorithm 3 in the Supplementary Material of Huang et al. (2025)) can be used for finding EW D-optimal designs, with $f({\mathbf{w}})$ replaced by $f_{\rm EW}({\mathbf{w}})$ or $f_{\rm SEW}({\mathbf{w}})$, called EW lift-one algorithm.
To speed up the lift-one algorithm, Yang and Mandal (2015) provided analytic solutions in their Lemma 4.2 for generalized linear models. Bu et al. (2020) mentioned but with no detail that the corresponding optimization problems for multinomial logistic models (MLM) also have analytic solutions when the number of response categories $J\leq 5$. In Section S1 of the Supplementary Material, we provide all the relevant formulae for the lift-one algorithm under an MLM with $J\leq 5$.
3.2 EW ForLion algorithm for EW D-optimal designs with mixed factors and a rounding algorithm for obtaining exact designs
When at least one factor is continuous, that is, at least $I_{1}=[a_{1},b_{1}]$, we follow the ForLion algorithm proposed by Huang et al. (2024) for locally D-optimal designs, to find EW D-optimal designs, called the EW ForLion algorithm (see Algorithm 1).
When there is no confusion, we may denote (i) ${\mathbf{F}}(\boldsymbol{\xi})$ for $E\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\Theta})\}$ under integral-based EW D-optimality or $\hat{E}\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\Theta})\}$ under sample-based EW D-optimality, for each $\boldsymbol{\xi}\in\boldsymbol{\Xi}$; and (ii) ${\mathbf{F}}_{\mathbf{x}}$ for $E\{{\mathbf{F}}({\mathbf{x}},\boldsymbol{\Theta})\}$ under integral-based EW D-optimality or $\hat{E}\{{\mathbf{F}}({\mathbf{x}},\boldsymbol{\Theta})\}$ under sample-based EW D-optimality, for each ${\mathbf{x}}\in{\cal X}$. As a result, $d({\mathbf{x}},\boldsymbol{\xi})$ in Theorem 2.1 can be simplified to $d({\mathbf{x}},\boldsymbol{\xi})={\rm tr}({\mathbf{F}}(\boldsymbol{\xi})^{-1}{\mathbf{F}}_{\mathbf{x}})$ (see Section S3 in the Supplementary Material on relevant efficient calculations).
@@ 3.1 Lift-one algorithm for EW D-optimal designs with discrete factors only
Step 3: Given $\boldsymbol{\xi}_{t}\in\boldsymbol{\Xi}$ after the merging step, apply the lift-one algorithm with the convergence threshold $\epsilon$, and obtain an EW D-optimal allocation $(w_{1}^{*},\ldots,w_{m_{t}}^{*})^{T}$ for the design points $\{{\mathbf{x}}_{1}^{(t)},\ldots,{\mathbf{x}}_{m_{t}}^{(t)}\}$. Replace $w_{i}^{(t)}$ with $w_{i}^{*}$ , respectively.
Step 4: Deleting step: Update $\boldsymbol{\xi}_{t}$ by discarding any ${\mathbf{x}}_{i}^{(t)}$ for which $w_{i}^{(t)}=0$.
Step 5: Adding new point: Given $\boldsymbol{\xi}_{t}\in\boldsymbol{\Xi}$, find $\mathbf{x}^{*}=(x_{1}^{*},\ldots,x_{d}^{*})^{T}\in{\cal X}$ that maximizes $d(\mathbf{x},\boldsymbol{\xi}_{t})$. More specifically, if all factors are continuous, $\mathbf{x}^{*}$ can be obtained by the “L-BFGS-B” quasi-Newton method (Byrd et al., 1995) directly. If the first $k$ factors are continuous with $11\leq k\leq d-$ 1, then we may first find
@@
for each $\mathbf{x}_{(2)}=(x_{k+1},\ldots,x_{d})^{T}\in\prod_{j=k+1}^{d}I_{j}$ , and then choose $\mathbf{x}_{(2)}^{*}$ that yields the largest $d(((\mathbf{x}_{(1)}^{*})^{T},\,\mathbf{x}_{(2)}^{T})^{T},\,\boldsymbol{\xi}_{t})$. Note that $\mathbf{x}_{(1)}^{*}$ depends on $\mathbf{x}_{(2)}$ Then $\mathbf{x}^{*}=((\mathbf{x}_{(1)}^{*})^{T},\,(\mathbf{x}_{(2)}^{*})^{T})^{T}$.
@@
Note that the EW ForLion algorithm is essentially applying the original ForLion algorithm proposed by Huang et al. (2024) to ${\mathbf{F}}(\boldsymbol{\xi})$ under an EW D-optimality. It should be noted that Step 2 in Algorithm 1 is different from the merging step in the original ForLion algorithm by addressing a possible issue when the merging step leads to a degenerated ${\mathbf{F}}(\boldsymbol{\xi}_{\rm mer})$, which is possible in practice.
According to Step 6 in Algorithm 1, $\boldsymbol{\xi}_{t}$ reported by the EW ForLion algorithm satisfies $\max_{{\mathbf{x}}\in{\cal X}}d({\mathbf{x}},\boldsymbol{\xi}_{t})\leq p$. As a direct conclusion of Theorem 2.1, we have the following corollary.
Corollary 3.1.
@@
Since the design found by Algorithm 1 is an approximate design with continuous design points, to facilitate practical uses, we propose a rounding algorithm (see Algorithm 2) to convert an approximate design to an exact design with user-specified set of grid points.
@@
Step 0: Input: An approximate design $\boldsymbol{\xi}=\{({\mathbf{x}}_{i},w_{i}),i=1,\ldots,m_{0}\}\in\boldsymbol{\Xi}$ with all $w_{i}>0$ and $|{\mathbf{F}}(\boldsymbol{\xi})|>0$, a pre-specified merging threshold $\delta_{r}$ (e.g., $\delta_{r}=0.1$, typically larger than $\delta$ in Algorithm 1), grid levels (or pace lengths) $L_{1},\ldots,L_{k}$ for the $k$ continuous factors (e.g., $L_{1}=\cdots=L_{k}=0.25$), respectively, and the total number $n$ of experimental units.
Step 1: Merging step: For each pair of $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$ , define their distance $d_{ij}=\left\|\mathbf{x}_{i}-\mathbf{x}_{j}\right\|$ if their levels of discrete factors are identical; and $\infty$ otherwise. If $d_{ij}<\delta_{r}$ , tentatively merge ${\mathbf{x}}_{1}$ and ${\mathbf{x}}_{2}$ into a single design point $(w_{i}\mathbf{x}_{i}+w_{j}\mathbf{x}_{j})/(w_{i}+w_{j})$ with combined weight $w_{i}+w_{j}$ , and denote the resulting design as $\boldsymbol{\xi}_{\rm mer}$ If $|{\mathbf{F}}(\boldsymbol{\xi}_{\rm mer})|>0$, replace $\boldsymbol{\xi}$ with $\boldsymbol{\xi}_{\rm mer}$; otherwise, do not merge. Stop if all such pairs have been examined.
@@
To implement the EW ForLion algorithm (Algorithm 1) for EW D-optimal designs, it is critical to calculate $E\{{\mathbf{F}}({\mathbf{x}},\boldsymbol{\Theta})\}$ or $\hat{E}\{{\mathbf{F}}({\mathbf{x}},\boldsymbol{\Theta})\}$ efficiently.
Example 3.1.
@@
Multinomial logistic models (MLM):For MLMs (Glonek and McCullagh, 1995; Zocchi and Atkinson, 1999; Bu et al., 2020), according to Theorem 2 of Huang et al. (2024)), it is enough to replace $u_{st}^{\mathbf{x}}$ , or $u_{st}^{\mathbf{x}}(\boldsymbol{\theta})$ as it is a function of $\boldsymbol{\theta}$, with $E\{u_{st}^{\mathbf{x}}(\boldsymbol{\Theta})\}=\int_{\boldsymbol{\Theta}}u_{st}^{\mathbf{x}}(\boldsymbol{\theta})Q(d\boldsymbol{\theta})$ or $\hat{E}\{u_{st}^{\mathbf{x}}(\boldsymbol{\Theta})\}=B^{-1}\sum_{j=1}^{B}u_{st}^{\mathbf{x}}(\hat{\boldsymbol{\theta}}_{j})$, for $s,t=1,\ldots,J$. The formulae for calculating $u_{st}^{\mathbf{x}}$ as a function of $\boldsymbol{\theta}$ can be found in Appendix A in Huang et al. (2024).
An alternative way is to utilize Corollary 3.1 in Bu et al. (2020). In our notation, ${\mathbf{F}}({\mathbf{x}},\boldsymbol{\theta})={\mathbf{X}}_{\mathbf{x}}^{T}{\mathbf{U}}_{\mathbf{x}}(\boldsymbol{\theta}){\mathbf{X}}_{\mathbf{x}}$ , where

### v2

@@ 2.2 Characteristics of EW D-optimal designs
Since the design region ${\cal X}$ typically contains infinitely many design points, which is different in nature from the experiments considered by Yang et al. (2016, 2017) and Bu et al. (2020), both the theoretical properties and numerical algorithms for EW D-optimal designs here are far more difficult.
@@
In this section, to characterize EW D-optimal designs, we fix either a prior distribution $Q(\cdot)$ on $\boldsymbol{\Theta}$ for integral-based EW D-optimality, or a set of bootstrapped or sampled parameter vectors $\{\hat{\boldsymbol{\theta}}_{1},\ldots,\hat{\boldsymbol{\theta}}_{B}\}$ for sample-based EW D-optimality, so that either $E\left\{{\mathbf{F}}({\mathbf{x}}_{i},\boldsymbol{\Theta})\right\}$ is defined as in (2.2), or $\hat{E}\left\{{\mathbf{F}}({\mathbf{x}}_{i},\boldsymbol{\Theta})\right\}$ is defined as in (2.4). Similarly to Fedorov and Leonov (2014) and Huang et al. (2024), we list the assumptions needed for this paper.
  • (A1)
    The design region ${\cal X}\subset\mathbb{R}^{d}$ is compact.
  • (A2)
    For each $\boldsymbol{\theta}\in\boldsymbol{\Theta}$, the Fisher information ${\mathbf{F}}({\mathbf{x}},\boldsymbol{\theta})$ is element-wise continuous with respect to all continuous factors of ${\mathbf{x}}\in{\cal X}$.
  • (A3)
    For each $\boldsymbol{\theta}\in\boldsymbol{\Theta}$, the Fisher information $\mathbf{F}(\mathbf{x},\boldsymbol{\theta})=(F_{st}({\mathbf{x}},\boldsymbol{\theta}))_{s,t=1,\ldots,p}$ is element-wise continuous with respect to all continuous factors of ${\mathbf{x}}\in{\cal X}$, and there exists a nonnegative and integrable function $K(\boldsymbol{\theta})$ on $\boldsymbol{\Theta}$, such that, $\int_{\boldsymbol{\Theta}}K(\boldsymbol{\theta})Q(d\boldsymbol{\theta})<\infty$, and $|F_{st}(\mathbf{x},\boldsymbol{\theta})|\leq K(\boldsymbol{\theta})$, for all $s,t\in\{1,\ldots,p\}$, $\mathbf{x}\in\mathcal{X}$, and $\boldsymbol{\theta}\in\boldsymbol{\Theta}$.
In this paper, ${\cal X}$ is either $\prod_{j=1}^{d}I_{j}$ or $\prod_{j=1}^{k}I_{j}\ \times{\cal D}$. Both of them are bounded and closed in $\mathbb{R}^{d}$ and thus Assumption (A1) is satisfied.
As for Assumptions (A2) and (A3), (i) if $k=0$, that is, all factors are discrete, then both Assumption (A2) and the first half (i.e., the continuous statement) of Assumption (A3) are automatically satisfied; (ii) when $k\geq 1$, that is, there is at least one continuous factor, (A3) always implies (A2).
Following Fedorov and Leonov (2014), to characterize the sample-based and integral-based D-optimal EW designs, we extend the collection $\boldsymbol{\Xi}$ of designs (each design consists only of a finite number of design points) to $\boldsymbol{\Xi}({\cal X})$, which consists of all probability measures on ${\cal X}$. In other words, a design $\boldsymbol{\xi}\in\boldsymbol{\Xi}({\cal X})$ is also a probability measure on ${\cal X}$, and for each $\boldsymbol{\theta}\in\boldsymbol{\Theta}$,
Then for sample-based EW criterion and $\boldsymbol{\xi}\in\boldsymbol{\Xi}({\cal X})$, we denote
${\mathbf{F}}_{\rm SEW}(\boldsymbol{\xi})=\hat{E}\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\Theta})\}=\frac{1}{B}\sum_{j=1}^{B}{\mathbf{F}}(\boldsymbol{\xi},\hat{\boldsymbol{\theta}}_{j})=\int_{\cal X}\hat{E}\{{\mathbf{F}}({\mathbf{x}},\boldsymbol{\Theta})\}\boldsymbol{\xi}(d{\mathbf{x}})\ .$ (2.5)
${\mathbf{F}}_{\rm EW}(\boldsymbol{\xi})$ $=$ $E\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\Theta})\}=\int_{\boldsymbol{\Theta}}{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\theta})Q(d\boldsymbol{\theta})$ (2.6)
[…]
Furthermore, we denote the collections of Fisher information matrices, ${\cal F}_{\rm SEW}({\cal X})=\{{\mathbf{F}}_{\rm SEW}(\boldsymbol{\xi})\mid\boldsymbol{\xi}\in\boldsymbol{\Xi}({\cal X})\}$, and ${\cal F}_{\rm EW}({\cal X})=\{{\mathbf{F}}_{\rm EW}(\boldsymbol{\xi})\mid\boldsymbol{\xi}\in\boldsymbol{\Xi}({\cal X})\}$. As a summary of the arguments in Section S2, we have the following lemma.
Lemma 1.
Under Assumptions (A1) and (A2), ${\cal F}_{\rm SEW}({\cal X})$ is convex and compact; while under Assumptions (A1) and (A3), ${\cal F}_{\rm EW}({\cal X})$ is convex and compact.
@@
With the aid of Lemma 1 and Carathéodory’s Theorem (see, for example, Theorem 2.1.1 in Fedorov (1972)), we obtain the following theorem, whose proof, as well as proofs for other theorems, is relegated to Section S3 in the Supplementary Material.
Theorem 1.
To characterize D-optimality for EW designs, we define the objective function $\Psi({\mathbf{F}})=-\log|{\mathbf{F}}|$ for ${\mathbf{F}}\in{\cal F}_{\rm SEW}({\cal X})$ or ${\cal F}_{\rm EW}({\cal X})$. Then $f_{\rm SEW}(\boldsymbol{\xi})=\exp\{-\Psi(\hat{E}\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\Theta})\})\}$, and $f_{\rm EW}(\boldsymbol{\xi})=\exp\{-\Psi(E\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\Theta})\})\}$. Note that minimizing $\Psi({\mathbf{F}})$ for ${\mathbf{F}}\in{\cal F}_{\rm SEW}({\cal X})$ or ${\cal F}_{\rm EW}({\cal X})$ is equivalent to maximizing $f_{\rm SEW}(\boldsymbol{\xi})$ or $f_{\rm EW}(\boldsymbol{\xi})$ for $\boldsymbol{\xi}\in\boldsymbol{\Xi}$, respectively, due to Theorem 1.
For D-optimality, according to Section 2.4.2 in Fedorov and Leonov (2014), $\Psi$ always satisfies their Assumptions (B1), (B2), and (B4). In our notations, we let $\boldsymbol{\Xi}(q)$ denote $\{\boldsymbol{\xi}\in\boldsymbol{\Xi}\mid|\hat{E}\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\Theta})\}|\geq q\}$ for sample-based EW D-optimality, or $\{\boldsymbol{\xi}\in\boldsymbol{\Xi}\mid|E\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\Theta})\}|\geq q\}$ for integral-based EW D-optimality. We still need the following assumption:
  • (B3)
    There exists a $q>0$, such that $\boldsymbol{\Xi}(q)$ is non-empty.
Theorem 2.
For sample-based EW D-optimality under Assumptions (A1), (A2), and (B3), or integral-based EW D-optimality under Assumptions (A1), (A3), and (B3), we must have (i) there exists an optimal design $\boldsymbol{\xi}^{*}$ that contains no more than $p(p+1)/2$ design points; (ii) the set of optimal designs is convex; and (iii) a design $\boldsymbol{\xi}$ is EW D-optimal if and only if $\max_{\mathbf{x}\in\mathcal{X}}d(\mathbf{x},\boldsymbol{\xi})\leq p$, where $d(\mathbf{x},\boldsymbol{\xi})={\rm tr}([E\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\Theta})\}]^{-1}E\{{\mathbf{F}}({\mathbf{x}},\boldsymbol{\Theta})\})$ for integral-based EW D-optimality, or ${\rm tr}([\hat{E}\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\Theta})\}]^{-1}\hat{E}\{{\mathbf{F}}({\mathbf{x}},\boldsymbol{\Theta})\})$ for sample-based EW D-optimality.
@@ 3 Algorithms for EW D-optimal Designs
3.1 EW lift-one algorithm for discrete factors only
When all factors are discrete, the design region ${\cal X}$ contains only a finite number of distinct experimental settings. In this case, we may denote the corresponding design settings as ${\mathbf{x}}_{1},\ldots,{\mathbf{x}}_{m}$ The EW D-optimal design problem in this case is to find ${\mathbf{w}}^{*}=(w_{1}^{*},\ldots,w_{m}^{*})^{T}\in S=\{(w_{1},\ldots,w_{m})^{T}\in\mathbb{R}^{m}\mid w_{i}\geq 0,i=1,\ldots,m;\sum_{i=1}^{m}w_{i}=1\}$, which maximizes $f_{\rm EW}({\mathbf{w}})=|\sum_{i=1}^{m}w_{i}E\{{\mathbf{F}}({\mathbf{x}}_{i},\boldsymbol{\Theta})\}|$ or $f_{\rm SEW}({\mathbf{w}})=|\sum_{i=1}^{m}w_{i}\hat{E}\{{\mathbf{F}}({\mathbf{x}}_{i},$ $\boldsymbol{\Theta})\}|$.
The lift-one algorithm (see Algorithm 3 in the Supplementary Material of Huang et al. (2025)) can be used for finding EW D-optimal designs, with $f({\mathbf{w}})$ replaced by $f_{\rm EW}({\mathbf{w}})$ or $f_{\rm SEW}({\mathbf{w}})$, called EW lift-one algorithm.
To speed up the lift-one algorithm, Yang and Mandal (2015) provided analytic solutions in their Lemma 4.2 for generalized linear models. Bu et al. (2020) mentioned but with no details that the corresponding optimization problems for multinomial logistic models (MLM) also have analytic solutions when the number of response categories $J\leq 5$. In Section S1 of the Supplementary Material, we provide all the relevant formulae for the lift-one algorithm under an MLM with $J\leq 5$.
3.2 EW ForLion algorithm and rounding algorithm
When at least one factor is continuous, we follow the ForLion algorithm proposed by Huang et al. (2024) for locally D-optimal designs, to find EW D-optimal designs, called the EW ForLion algorithm (see Algorithm 1).
When there is no confusion, we denote (i) ${\mathbf{F}}(\boldsymbol{\xi})$ for $E\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\Theta})\}$ under integral-based EW D-optimality or $\hat{E}\{{\mathbf{F}}(\boldsymbol{\xi},\boldsymbol{\Theta})\}$ under sample-based EW D-optimality, for each $\boldsymbol{\xi}\in\boldsymbol{\Xi}$; and (ii) ${\mathbf{F}}_{\mathbf{x}}$ for $E\{{\mathbf{F}}({\mathbf{x}},\boldsymbol{\Theta})\}$ under integral-based EW D-optimality or $\hat{E}\{{\mathbf{F}}({\mathbf{x}},\boldsymbol{\Theta})\}$ under sample-based EW D-optimality, for each ${\mathbf{x}}\in{\cal X}$. As a result, $d({\mathbf{x}},\boldsymbol{\xi})$ in Theorem 2 can be simplified to $d({\mathbf{x}},\boldsymbol{\xi})={\rm tr}({\mathbf{F}}(\boldsymbol{\xi})^{-1}{\mathbf{F}}_{\mathbf{x}})$.
@@ 3.2 EW ForLion algorithm and rounding algorithm
Step 3: Given $\boldsymbol{\xi}_{t}\in\boldsymbol{\Xi}$ after the merging step, apply the lift-one algorithm (see Algorithm 3 in the Supplementary Material of Huang et al. (2025) and Remark 1 in Huang et al. (2024)) with the convergence threshold $\epsilon$, and obtain an EW D-optimal allocation $(w_{1}^{*},\ldots,w_{m_{t}}^{*})^{T}$ for the design points $\{{\mathbf{x}}_{1}^{(t)},\ldots,{\mathbf{x}}_{m_{t}}^{(t)}\}$. Replace $w_{i}^{(t)}$ with $w_{i}^{*}$ , respectively.
Step 4: Deleting step: Update $\boldsymbol{\xi}_{t}$ by discarding any ${\mathbf{x}}_{i}^{(t)}$ with $w_{i}^{(t)}=0$.
Step 5: Adding new point: Given $\boldsymbol{\xi}_{t}\in\boldsymbol{\Xi}$, find $\mathbf{x}^{*}\in{\cal X}$ that maximizes $d(\mathbf{x},\boldsymbol{\xi}_{t})$. More specifically, if all factors are continuous, $\mathbf{x}^{*}$ can be obtained by the “L-BFGS-B” quasi-Newton method (Byrd et al., 1995) directly. If the first $k$ factors are continuous with $11\leq k\leq d-$ 1, we first find
@@
for each $\mathbf{x}_{(2)}\in{\cal D}\subseteq\prod_{j=k+1}^{d}I_{j}$ , and then choose $\mathbf{x}_{(2)}^{*}$ that yields the largest $d(((\mathbf{x}_{(1)}^{*})^{T},\,\mathbf{x}_{(2)}^{T})^{T},\,\boldsymbol{\xi}_{t})$. Note that $\mathbf{x}_{(1)}^{*}$ depends on $\mathbf{x}_{(2)}$ Then $\mathbf{x}^{*}=((\mathbf{x}_{(1)}^{*})^{T},\,(\mathbf{x}_{(2)}^{*})^{T})^{T}$.
@@
The EW ForLion algorithm is essentially applying the original ForLion algorithm proposed by Huang et al. (2024) to ${\mathbf{F}}(\boldsymbol{\xi})$ under an EW D-optimality. It should be noted that (i) Step 2 in Algorithm 1 is different from the merging step in the original ForLion algorithm by addressing a possible issue when the merging step leads to a degenerated ${\mathbf{F}}(\boldsymbol{\xi}_{\rm mer})$, which is possible in practice; and (ii) in Step 5, a general subset ${\cal D}$ is used instead of $\prod_{j=k+1}^{d}I_{j}$ , which is also different from the original ForLion algorithm (see Examples 1 and 2 for motivations).
According to Step 6 in Algorithm 1, $\boldsymbol{\xi}_{t}$ reported by the EW ForLion algorithm satisfies $\max_{{\mathbf{x}}\in{\cal X}}d({\mathbf{x}},\boldsymbol{\xi}_{t})\leq p$. As a direct conclusion of Theorem 2, we have the following corollary.
Corollary 1.
@@
Similarly to the ForLion algorithm (Huang et al., 2024), we may relax in practice the stopping rule $d\left(\mathbf{x}^{*},\boldsymbol{\xi}_{t}\right)\leq p$ in Step 6 to $d\left(\mathbf{x}^{*},\boldsymbol{\xi}_{t}\right)\leq p+\epsilon$. Following the arguments in Remark 2 in Huang et al. (2024), as a direct conclusion of Theorem 2, Algorithm 1 is guaranteed to stop in finite steps.
Since the design found by Algorithm 1 is an approximate design with continuous design points, to facilitate users in practice, we propose a rounding algorithm (see Algorithm 2) to convert an approximate design to an exact design with a user-specified set of grid points. Different from the rounding algorithms proposed in the literature (see, for example, Algorithm 2 in Huang et al. (2025)), Algorithm 2 here involves rounding both the weights and the design points.
@@
Step 0: Input: An approximate design $\boldsymbol{\xi}=\{({\mathbf{x}}_{i},w_{i}),i=1,\ldots,m_{0}\}\in\boldsymbol{\Xi}$ with all $w_{i}>0$ and $|{\mathbf{F}}(\boldsymbol{\xi})|>0$, a pre-specified merging threshold $\delta_{r}$ (e.g., $\delta_{r}=0.1$, typically larger than $\delta$ in Algorithm 1), grid levels (or pace lengths) $L_{1},\ldots,L_{k}$ for the $k$ continuous factors (e.g., $L_{1}=\cdots=L_{k}=0.25$, typically larger than $\delta_{r}$), respectively, and the total number $n$ of experimental units.
Step 1: Merging step: For each pair of $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$ , define their distance $d_{ij}=\left\|\mathbf{x}_{i}-\mathbf{x}_{j}\right\|$ if their levels of discrete factors are identical; and $\infty$ otherwise. If $d_{ij}<\delta_{r}$ , tentatively merge ${\mathbf{x}}_{i}$ and ${\mathbf{x}}_{j}$ into a single design point $(w_{i}\mathbf{x}_{i}+w_{j}\mathbf{x}_{j})/(w_{i}+w_{j})$ with combined weight $w_{i}+w_{j}$ , and denote the resulting design as $\boldsymbol{\xi}_{\rm mer}$ If $|{\mathbf{F}}(\boldsymbol{\xi}_{\rm mer})|>0$, replace $\boldsymbol{\xi}$ with $\boldsymbol{\xi}_{\rm mer}$; otherwise, do not merge. Stop if all such pairs have been examined.
@@
As illustrated by Yang et al. (2016, 2017) and Bu et al. (2020), one major advantage of EW D-optimality against Bayesian D-optimality is that it optimizes $|E\{{\mathbf{F}}({\mathbf{x}},\boldsymbol{\Theta})\}|$ or $|\hat{E}\{{\mathbf{F}}({\mathbf{x}},\boldsymbol{\Theta})\}|$, instead of $E\{\log|{\mathbf{F}}({\mathbf{x}},\boldsymbol{\Theta})|\}$. To implement the EW ForLion algorithm (Algorithm 1) for EW D-optimal designs, in this section, we show by two examples that instead of calculating ${\mathbf{F}}_{\mathbf{x}}=E\{{\mathbf{F}}({\mathbf{x}},\boldsymbol{\Theta})\}$ or $\hat{E}\{{\mathbf{F}}({\mathbf{x}},\boldsymbol{\Theta})\}$ directly, we may focus on key components of the Fisher information matrix to make the computation more efficiently.
Example 3.
@@
Multinomial logistic models (MLM):A general MLM (Glonek and McCullagh, 1995; Zocchi and Atkinson, 1999; Bu et al., 2020) for categorical responses with $J$ categories can be defined by ${\mathbf{C}}^{T}\log({\mathbf{L}}\boldsymbol{\pi}_{i})=\boldsymbol{\eta}_{i}={\mathbf{X}}_{i}\boldsymbol{\theta}$, $i=1,\ldots,m$, with $(2J-1)\times J$ constant matrices ${\mathbf{C}}$ and ${\mathbf{L}}$, category probabilities $\boldsymbol{\pi}_{i}=(\pi_{i1},\ldots,\pi_{iJ})^{T}$, $J\times p$ model matrices ${\mathbf{X}}_{i}$ , and model parameters $\boldsymbol{\theta}\in\boldsymbol{\Theta}\subseteq\mathbb{R}^{p}$. According to Theorem 2 of Huang et al. (2024), to calculate the Fisher information ${\mathbf{F}}_{\mathbf{x}}$ at ${\mathbf{x}}\in{\cal X}$, it is enough to replace $u_{st}^{\mathbf{x}}$ , or $u_{st}^{\mathbf{x}}(\boldsymbol{\theta})$ as it is a function of $\boldsymbol{\theta}$, with $E\{u_{st}^{\mathbf{x}}(\boldsymbol{\Theta})\}=\int_{\boldsymbol{\Theta}}u_{st}^{\mathbf{x}}(\boldsymbol{\theta})Q(d\boldsymbol{\theta})$ or $\hat{E}\{u_{st}^{\mathbf{x}}(\boldsymbol{\Theta})\}=B^{-1}\sum_{j=1}^{B}u_{st}^{\mathbf{x}}(\hat{\boldsymbol{\theta}}_{j})$, for $s,t=1,\ldots,J$. The formulae for calculating $u_{st}^{\mathbf{x}}(\boldsymbol{\theta})$ can be found in Appendix A of Huang et al. (2024). Then ${\mathbf{F}}({\mathbf{x}},\boldsymbol{\theta})={\mathbf{X}}_{\mathbf{x}}^{T}{\mathbf{U}}_{\mathbf{x}}(\boldsymbol{\theta}){\mathbf{X}}_{\mathbf{x}}$ according to Corollary 3.1 in Bu et al. (2020), where